OpenStudy (anonymous):

Solve the following system using the substitution method. 3x - y = -2 6x + 5y = 24 negative two-thirds, negative four negative two-thirds, four two-thirds, negative four two-thirds, four

OpenStudy (anonymous):

@matricked

OpenStudy (anonymous):

can you isolate y from first eq 3x - y = -2

OpenStudy (anonymous):

y = 2 + 3x

OpenStudy (anonymous):

3x - y = -2 so we can rewrite 3x+2 =y or y=3x+2 now we can plug y=3x+2 in 6x + 5y = 24

OpenStudy (anonymous):

yup u r doing right ..

OpenStudy (anonymous):

now in 6x + 5y = 24 replace y by 2 + 3x

OpenStudy (anonymous):

6x + 5 (3x + 2) = 24 6x + 15x +10 = 24 21x = 14

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

14/21 = 2/3

OpenStudy (anonymous):

so thats x

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

now use x =2/3 in y=2+3x to find y

OpenStudy (anonymous):

and its -4 so the answer is C

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

are you getting y as -4 y=2+3*(2/3) =2+2 =+4

OpenStudy (anonymous):

so D

OpenStudy (anonymous):

oh thanks